OpenStudy (anonymous):

Simplify 1/5^-4 -4/5 625 -625 -1/625

OpenStudy (anonymous):

You're aware of laws of exponents by ANY chance?

OpenStudy (anonymous):

http://gradeninemath.blogspot.in/2010/06/number-sense.html for laws of exponents.

OpenStudy (anonymous):

so it says that a^-n = 1/a^n. u can apply conversely also. a^n = 1/a^-n. u can try out the given sum now?

OpenStudy (anonymous):

-625?

OpenStudy (anonymous):

actually nope. bcoz 1/5^-4 would be 5^4 which is 625.